Firewall Project XT. Using the “complexity weighting” scheme and the function point complexity weighted table shown below, calculate the total function point count. Assume historical data suggest 6 function points equal one person-month and 7 people can work on the project. (Round the final answers to the nearest whole numbers.)

Complexity Weight Table
Number of inputs 8 Rated complexity low 3 points
Number of outputs 19 Rated complexity average 5 points
Number of inquiries 10 Rated complexity average 5 points
Number of files 35 Rated complexity high 11 points
Number of interfaces 60 Rated complexity high 13 points


a. The total function point count is:



b. Calculate the estimated project duration:



c. If 25 people are available for the project, what is the estimated project duration?



d. If the project must be complete in six months, how many people will be needed for the project?
Answer #1

a.

Total function point count = 1334

Element Count Low Average High Total
Inputs 8 x 3 24
Outputs 19 x 5 95
Inquiries 10 x 5 50
Files 35 x 11 385
Interfaces 60 x 13 780
Total 1334

b.

6 points = 1 person-month

So, no. of person-month required = 1334 / 6 = 222.333

Given that 7 people can work on the project, the expected duration = 222.333 / 7 = 32 months

c.

Given that 25 people can work on the project, the expected duration = 222.333 / 25 = 9 months

d.

Person-month = 222.33

Duration in months = 6

So, no. of persons required = 222.33 / 6 = 37